The specific heat of a 90 g block of material is to be determined. The block is placed in a 25 g copper calorimeter that also holds 60 g of water. The system is initially at 20
Answer #1

\sum Q = 0

m_1c_1\Delta T_1 +m_2c_2\Delta T_2 +m_3c_3\Delta T_3+m_4c_4\Delta T_4 =0

90c_1(54-20)+25(0.092)(54-20)+60(1)(54-20)+117(1)(54-80) = 0

solving for c1 we have:

c_1 = 0.302 cal/g C

solving for c1 we get:

c_1 = 0.909 ca/g C